GENERALIZED BERGMANN METRICS AND INVARIANCE OF PLURIGENERA

HAJIME TSUJI

Abstract. An invariant kernel for the pluricanonical system of a projective manifold of general type is introduced. Using this kernel we prove that the Yau volume form on a smooth projective variety has seminegative Ricci curvature. As a biproduct we prove the invariance of plurigenera for smooth projective deformations of manifolds of general type.